OpenStudy (anonymous):

Help please..what is the vertex form of the equation? y=x^2+12x-4 Pleasee help!

OpenStudy (anonymous):

a(x – h)^2 + k is vertex form

OpenStudy (anonymous):

what

OpenStudy (anonymous):

I can't see what that says

OpenStudy (anonymous):

a(x – h)^2 + k

OpenStudy (anonymous):

a(x - h)^2 + k

OpenStudy (anonymous):

Where h,k is the vertex

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

factors of -4 that add up to 12 are -4 and 16

OpenStudy (anonymous):

y = x^2 + 12x - 4 y = x^2 + 12x + (36 - 36) - 4

OpenStudy (anonymous):

y = (x^2 + 12x + 36) - 36 - 4 y = (x + 6)^2 - 40

OpenStudy (anonymous):

@swagmaster47

OpenStudy (anonymous):

@mathmate is this right what I did?

OpenStudy (mathmate):

What you did is right, but not yet complete!

OpenStudy (mathmate):

You need the vertex, which is (h,k) in \(y=(x-h)^2+k\) Interpret the values of h and k and you're done!

OpenStudy (anonymous):

ohh ok but wait I just realized that there was a negative sign in front of the x

OpenStudy (mathmate):

Exactly, you have just avoided a very common mistake, good job. So what do you find for the vertex (h,k) ?

OpenStudy (anonymous):

y = -x^2 + 12x - 4 y= - [ (x^2 -12x +36)- 36 +4] y= - ( x- 6)^2 -32 ( x- 6)^2 = - y-32

OpenStudy (mathmate):

\(y= - [ (x^2 -12x +36)- 36 \color{red}{-}4] \) What you had was correct, just need to interpret h and k, from \(y = (x + 6)^2 - 40 \) and \(y=(x-h)^2+k\) The vertex is at (h,k).

OpenStudy (mathmate):

there is no initial negative in \(y= [ (x^2 -12x +36)- 36 \color{red}{-}4]\)

OpenStudy (anonymous):

oh ok

OpenStudy (mathmate):

So what do you get for h and k?

OpenStudy (mathmate):

Compare the two equations which are equivalent, then try to deduce the value of h and k: \(y = (x + 6)^2 - 40 ~~\equiv~~ y=(x-h)^2+k\)

OpenStudy (anonymous):

Now it works

OpenStudy (mathmate):

Do you see better this way? I have a problem with computer loading as well, it's very slow, don't know if it's me or the server.

OpenStudy (anonymous):

Yes I do! :) thank you

OpenStudy (mathmate):

You're welcome! :) BTW, what are the values you found for h & k?

OpenStudy (anonymous):

they were -8 and 4

OpenStudy (mathmate):

Wait... (x+6)=(x-h) so what is h?

OpenStudy (anonymous):

4

OpenStudy (mathmate):

:( \(x+6\ne x-4\)

OpenStudy (anonymous):

6? Idk sorry :(

OpenStudy (mathmate):

\(x+6\ne x-6\) You need to do something!

OpenStudy (anonymous):

I see is it 12>

OpenStudy (mathmate):

\(x+6 = x-(-6)\), so h=-6 is that ok?

OpenStudy (anonymous):

Yes

OpenStudy (mathmate):

Now you need to find k! Compare both sides of the equation, find k so that both sides will be equal. \(y = (x + 6)^2 - 40 ~~\equiv~~ y=(x-h)^2+k\)

OpenStudy (mathmate):

The server or my computer is very slow. I will take a break!